OpenStudy (anonymous):

How to integrate 3^(x+3^x)

sam (.sam.):

Try U-sub

OpenStudy (anonymous):

Okay sure

OpenStudy (welshfella):

splitting it up first might help 3^(x+3^x) = 3^x * 3^(3^x) = 3^x * 3^(3x)

OpenStudy (anonymous):

So the integral of 3^x is 3^x / ln(3)

sam (.sam.):

\[\int\limits 3^{x+3^x}dx = \int\limits 3^x \times 3^{3^x} dx\] Let \[u=3^x \\ \\ du=3^xln(3)dx\] \[\int\limits 3^u\frac{1}{\ln(3)}du\]

sam (.sam.):

Continue

OpenStudy (rational):

@welshfella \[\large 3^{(3^x)}~~ \ne~~ 3^{3x}\] right

OpenStudy (anonymous):

Oh yes! I got it! The answer is 3^(3^x) / (ln(3))^2 + C

OpenStudy (anonymous):

So when I look at these types of problems I should look at splitting up the problem and if I can u-sub something. Thanks

OpenStudy (anonymous):

Thank you everyone~

sam (.sam.):

Yes always think of U-sub first, then followed by other methods if you can't U-sub

OpenStudy (welshfella):

@rational - right!